# The combinatorics of some two-color partition identities

### v1

@@ 1. Introduction
An overpartition of $n$ is a partition of $n$ where the first occurrence of each part may be overlined and denote the cardinality of it by $\overline{p}(n)$. For example, $(\overline{5},5,3,2,\overline{1},1)$ is an overpartition of 1717. We denote the number of overpartitions of $n$ into odd parts by $\overline{p}_{o}(n)$.
@@
Theorem $1.$ ([9] Theorem 1.3.) For any nonnegative integer $n$,
@@ 3. The proof of $(f)$ and $(g)$
Before providing a combinatorial proof for $(f)$ and $(g)$, we first introduce the 44-modular diagram of $\lambda$ which all parts are distinct and the even part must be a multiple of 44. For $\lambda$, we can decompose it into three parts, $\lambda_{e}$ which parts are even, $\lambda_{C1}$ which parts are congruent to 11 modulo 44, and $\lambda_{C3}$ which parts are congruent to 33 modulo 44. In the 44-modular diagram, $\lambda_{C3}$ parts are represented in rows above the main diagonal and $\lambda_{C1}$ parts are represented in columns below the main diagonal ( $\lambda_{e}$ parts are not represented in the diagram). For $4k+1$, its representation would have $k$ squares and 11 triangle and write 11 inside the triangle. For $4k+3$, its representation would have $k$ squares and 11 triangle and write 33 inside the triangle.
@@
CombinatorialCombinatorial proofproof forfor $(f)$ andand $(g)$. For $(f)$, we let $\pi_{L}(n)$ the set of two-colour partitions of $n$ in which all parts are distinct and the odd parts may occur only in the blue colour. We also denote $\pi_{M}(n)$ the subset of $\pi_{L}(n)$ in which there exists at least an even number in only one color. Also, we let $\pi_{N}(n)$ the set of partitions of $n$ in which all parts are distinct and the even part must be a multiple of 44.
@@
Example $3.$ Let us give an example.

### v2

@@ 1. Introduction
An overpartition of $n$ is a partition of $n$ where the first occurrence of each part may be overlined and denote the cardinality of it by $\overline{p}(n)$. For example, $(\overline{5},5,3,2,\overline{1},1)$ is an overpartition of $17$. We denote the number of overpartitions of $n$ into odd parts by $\overline{p}_{o}(n)$.
@@
Theorem 1.1. ([9] Theorem 1.3.) For any nonnegative integer $n$,
@@ 3. The proof of $(f)$ and $(g)$
Before providing a combinatorial proof for $(f)$ and $(g)$, we first introduce the $4$-modular diagram of $\lambda$ which all parts are distinct and the even part must be a multiple of $4$. For $\lambda$, we can decompose it into three parts, $\lambda_{e}$ which parts are even, $\lambda_{C1}$ which parts are congruent to $1$ modulo $4$, and $\lambda_{C3}$ which parts are congruent to $3$ modulo $4$. In the $4$-modular diagram, $\lambda_{C3}$ parts are represented in rows above the main diagonal and $\lambda_{C1}$ parts are represented in columns below the main diagonal ( $\lambda_{e}$ parts are not represented in the diagram). For $4k+1$, its representation would have $k$ squares and $1$ triangle and write $1$ inside the triangle. For $4k+3$, its representation would have $k$ squares and $1$ triangle and write $3$ inside the triangle.
@@
CombinatorialCombinatorial proofproof forfor $(f)$ andand $(g)$. For $(f)$, we let $\pi_{L}(n)$ the set of two-colour partitions of $n$ in which all parts are distinct and the odd parts may occur only in the blue colour. We also denote $\pi_{M}(n)$ the subset of $\pi_{L}(n)$ in which there exists at least an even number in only one color. Also, we let $\pi_{N}(n)$ the set of partitions of $n$ in which all parts are distinct and the even part must be a multiple of $4$.
@@
Example 3.3. Let us give an example.